# A parameterized block-splitting preconditioner for indefinite least squares problem

### v1

@@
Abstract. We present a stationary iteration based upon a block splitting for a class of indefinite least squares problem. Convergence of the proposed method is investigated and optimal value of the involving parameter is used. The induced preconditioner is applied to accelerate the convergence of the GMRES method for solving the problem. We also analysed the eigenpair distribution of the preconditioned matrix. Some numerical are presented to show the effectiveness of the preconditioner. Numerical comparison with other well-known methods are also presented.
Keywords: Block, splitting, least squares, indefinite, preconditioning, iterative.
AMS Subject Classification
: 65F10, 65F50, 65F08..
@@ 1 Introduction
$Q=\begin{pmatrix}I_{p}&0\\ 0&I_{q}\end{pmatrix},\qquad p+q=m.$ (2)
@@
The normal equations corresponding to the problem (1) can be written as
@@
It follows from (3) that unlike the least squares problem, the ILS problem does not necessarily have a solution. The Hessian matrix for the function $f(x)=(b-Ax)^{T}J(b-Ax)$ is $H=A^{T}JA$. So the minimization problem (1) has a unique solution if and only if the matrix $A^{T}JA$ is symmetric positive definite (SPD) (see [4, 7]). When the problem size is small, Chandrasekaran et al. in [4] proposed a direct method based on the QR decomposition and Cholesky factorization for solving the problem. However, when the problem is large, the computational cost becomes prohibitively high. So it is recommended to use iterative methods for solving the problem.
In [9], Liu and Liu presented the two versions of the block SOR method for solving the problem. In [15], Song constructed an unsymmetric SOR (USSOR) method for solving the problem. A variable parameter Uzawa method for solving the indefinite least squares problem was presented by Meng et al. in [12]. Meng et al. in [11] proposed an enhanced Landweber method with momentum acceleration for solving large indefinite least squares problem. Recently, Meng et al. in [10] have stated alternating direction implicit iterative method for the problem.
@@
$A=\begin{pmatrix}A_{1}\cr A_{2}\end{pmatrix},\qquad A=\begin{pmatrix}b_{1}\cr b_{2}\end{pmatrix}$ (4)
@@
$\mathcal{B}$ $=$ $\begin{pmatrix}I&0&0\cr 0&P&0\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&-A_{1}&0\cr 0&0&-A_{2}^{T}\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{1}-\mathcal{N}_{1},$ (6)
$\mathcal{B}$ $=$ $\begin{pmatrix}I&0&0\cr 0&P&A_{2}^{T}\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&-A_{1}&0\cr 0&0&0\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{2}-\mathcal{N}_{2},$ (7)
$\mathcal{B}$ $=$ $\begin{pmatrix}I&A_{1}&0\cr 0&P&0\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&0&0\cr 0&0&-A_{2}^{T}\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{3}-\mathcal{N}_{3},$ (8)
and proved that the stationary iteration methods corresponding to the above splittings are unconditionally convergent. They applied the matrices $\mathcal{M}_{i}$, $i=1,2,3$ as preconditioners to accelerate the convergence of GMRES for solving the system (9). In this paper, we propose a parameterized block splitting for another formulation of the equation (3) used in [15], investigate convergence of the method and induced preconditioner.
@@ 2 The proposed iteration method
We consider the following reformulation of the problem (3) used in [15]
where $I$ denotes the identity matrix of the appropriate dimensions and $\delta=(\delta_{1};\delta_{2})=b-Ax$. It is noted that the coefficient matrix of this system is $\textbf{n}=2n+q$. By premultiplying the first row of the above system and letting $\hat{\delta}_{1}=A^{T}\delta_{1}$ and $\hat{b}_{1}=A_{1}^{T}b$, we obtain the following system
where $P=A_{1}^{T}A_{1}$. In this case, all the block diagonal matrices are nonsingular. We now consider the stationary iterative method induced by the parametrized block splitting (PBS) $\mathcal{A}=\mathcal{M}_{\alpha}-\mathcal{N}_{\alpha}$, where
@@
$\begin{cases}x^{(k+1)}=-P^{-1}(\hat{\delta}_{1}^{(k)}+\hat{b}_{1}),\\ \delta_{2}^{(k+1)}=-\alpha A_{2}x^{(k+1)}+(\alpha-1)A_{2}x^{(k)}+b_{2},\qquad k=0,1,\ldots,\\ \hat{\delta}_{1}^{(k+1)}=A_{2}^{T}\delta_{2}^{(k+1)},\end{cases}$ (13)
@@
The iterative scheme (12) converges to the solution of the system (10) for any arbitrary initial guess if and only if $\rho(\mathcal{G}_{\alpha})<1$ [13]. In the sequel we investigate the convergence of the proposed method. We first recall two lemmas.
@@
Let $A\in\mathbb{R}^{m\times n}$ and the matrix $A^{T}JA$ be SPD, where the matrix $J$ was defined in Eq. (2). Suppose that the matrix $A_{1}$ defined in Eq. (4) is of full column rank. If $\mu$ is an eigenvalue of the matrix $Q=(A_{1}^{T}A_{1})^{-1}(A_{2}^{T}A_{2})$, then $10\leq\mu<$ 1.
@@
It follows from Lemma 3.2 in [9] or Theorem 7.7.3 in [8]. ∎
@@
Let $(\lambda,\textbf{x})$ be an eigenpair of the iteration matrix $\mathcal{G}_{\alpha}$, i.e., $\mathcal{G}_{\alpha}\textbf{x}=\lambda\textbf{x}$. This means that $\mathcal{N}_{\alpha}\textbf{x}=\lambda\mathcal{M}_{\alpha}\textbf{x}$. So if we assume that $\textbf{x}=(x;y;z)$, then
@@
From Lemma 2, we have $10\leq\mu<$ 1. For $\mu=0$, both of the above inequalities hold true. So, we assume that $\mu\neq 0$. Clearly, the second equation holds true. From the first equation we see that
@@
$\alpha_{opt}={\rm arg}\min_{0<\alpha<1+\frac{1}{\mu_{\max}}}\rho(\mathcal{G}_{\alpha})=\frac{2}{1+\sqrt{1-\mu_{\max}}}.$
@@
We consider the equation (15) and without lose of generality we assume that $\mu=0$. In this case, this equation has two roots
@@ 3 The PBS preconditioner
In the previous section, the PBS method is convergent if and only if $(0,1+1/\mu_{\max})$. It means that the eigenvalues of the matrix
$\mathcal{M}_{\alpha}^{-1}\mathcal{A}=\mathcal{I}-\mathcal{G}_{\alpha},$
@@
$\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}z_{1}\cr z_{2}\cr z_{3}\end{pmatrix}=\begin{pmatrix}w_{1}\cr w_{2}\cr w_{3}\end{pmatrix},$
can be done via the following steps
@@
    Compute $z_{3}=w_{3}+A_{2}^{T}z_{2}$.
The coefficient matrix of the system in Step 1 is SPD. So, it can be solved directly using the Cholesky factorization for inexactly using the conjugate gradient method. The next theorem further analyses the eigenpairs distribution of the preconditioned matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$.
@@
    If $\lambda\neq 1$ (for every $\alpha$), then the corresponding eigenvector is given by
    $\left(\frac{1}{\lambda-1}P^{-1}A_{2}^{T}y;y;A_{2}^{T}y\right),$
@@
    If $\alpha=\lambda=1$, then every vector of the form $(x;y;0)$ with at least one of the vectors $x$ and $y$ being nonzero is the corresponding eigenvector.
@@
    If $\alpha\neq 1$ and $\lambda=1$, then every nonzero vector $(x;y;0)$ with $x\in\mathcal{N}(A_{2})$ is the corresponding eigenvector.
@@
Let $(\lambda,\textbf{x})$ be an eigenpair of the matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$, i.e., $\mathcal{M}_{\alpha}^{-1}\mathcal{A}\textbf{x}=\lambda\textbf{x}$. Letting $\textbf{x}=(x;y;z)$, this can be written as
$\begin{pmatrix}P&0&I\cr A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr y\cr z\end{pmatrix}=\lambda\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr y\cr z\end{pmatrix},$
@@
$\begin{cases}(\lambda-1)Px=z,\\ (1-\alpha\lambda)A_{2}x=(\lambda-1)y,\\ (\lambda-1)(z-A_{2}^{T}y)=0.\end{cases}$ (17)
To prove $(i)$, we first show that $y\neq 0$. If $y=0$, then from the third equation in (17) we get $z=0$. Substituting $z=0$ in the first equation yields $x=0$. So, in this case we deduce that $\textbf{x}=(x;y;z)=0$ which is a contradiction with the vector x being an eigenvector. Now, from the third equation we see that $z=A_{2}^{T}y$. Substituting $z=A_{2}^{T}y$ in the first equation
$x=\frac{1}{\lambda-1}P^{-1}z=\frac{1}{\lambda-1}P^{-1}A_{2}^{T}y.$
To prove $(ii)$, if $\alpha=\lambda=1$ then it follows from (17) that $z=0$. It is noted that for every vectors $x$ and $y$ the Eq. (17) holds true that proves $(ii)$.
To prove $(iii)$, we see that $z=0$ and $(1-\alpha)A_{2}x=0$. We conclude that $A_{2}x=0$, meaning that $x\in\mathcal{N}(A_{2})$. Hence the desired results is obtained. ∎
@@ 4 Numerical experiments
I this section we present some numerical experiments to verify the theoretical results and show the effectiveness of the PBS preconditioner. Comparison with other recently proposed preconditioners are also given. All the numerical experiments have been done by some Matlab codes with a Laptop with 2.40 GHz central processing unit (Intel(R) Core(TM) i7-5500), 8 GB memory and Windows 10 operating system.
@@
$A_{1}^{T}A_{1}-A_{2}^{T}A_{2}=\begin{pmatrix}35&10&16\\ 10&11&19\\ 16&19&44\\ \end{pmatrix},$
which is an SPD matrix. So all the conditions of Theorems 1 and 2 hold true. On the other hand, we have $\mu_{\max}=0.4976$. So, the method is convergent for any $\alpha$ with
$3.0090<\alpha<1+\frac{1}{\mu_{\max}}=1+\frac{1}{0.4976}=$ 3.009 .
@@
We have solved the system (10) using the PBS method (Eq. (14)). The initial guess is set to be a zeros vector and the iteration is stopped as soon as residual norm of the system is reduced by a factor of $10^{11}$. The number of iterations for the convergence for the optimum value $\alpha$ and some other values have been presented in Table 1. The numerical results confirm the theoretical results.
@@
In this example, we use four matrices TOLS340,TOLS1090, TOLS2000 and TOLS400 from Matrix Market (https://math.nist.gov/MatrixMarket/). The Tolosa matrices are encountered in the stability analysis of an aircraft flight model. The notable modes of this system are characterized by complex eigenvalues, with their imaginary parts situated within a specified frequency range. The objective is to determine the eigenvalues that possess the largest imaginary components. This problem has been studied at CERFACS (Centre Européen de Recherche et de Formation Avancée en Calcul Scientifique) in collaboration with the Aerospatiale Aircraft division. These matrices with their generic properties, including the size ($n$), the number of nonzero entries ($nnz$) and the condition number estimation (Cond), are presented in Table 2.
@@
matrix     $n$      $nnz$      Cond
TOLS 340 340 2196 $052.4e+$ 05
TOLS1090 1090 3546 $062.1e+$ 06
TOLS2000 2000 5184 $066.9e+$ 06
TOLS4000 4000 8784 $072.7e+$ 07
We set $A_{1}=$ TOLS 340, TOLS1090, TOLS2000 and TOLS4000. In this case, we have $p=n=340,1090,2000$ and $4000$. For each system we set $q=n$ and $A_{2}=6I_{n}$, where $I_{n}$ is the identity matrix of order $n$. The entries of the vectors $b_{1}$ and $b_{2}$ are set to be all ones. We solve the system (10) with restarted version of GMRES(10) in conjunction with the PBS preconditioner. Left-preconditioning is used. A zero vector is used as an initial guess and the iteration is stopped as soon as the residual of the original system is reduced by a factor of $10^{11}$. The parameter $\alpha$ in the PBS method is set to be $1$. We also compare the numerical results of PBS preconditioner with the preconditioners induced by the block splittings (6), (7) and (8) (denoted by BS1, BS2 and BS3) applied to the system (10). Numerical results have been presented in Table 2. In this table the number of iterations (Iter), elapsed CPU time (CPU),
@@
[…]
@@
   $m\times n$ PBS BS1 BS2 BS3 No-Prec
[…]
CPU 0.08 0.10 0.09 0.10
RES 9.13e-12 3.79e-12 3.52e-13 7.31e-12
ERR 9.22e-12 3.07e-11 9.16e-13 9.06e-12
[…]
CPU 0.09 0.11 0.10 0.11
RES 2.58e-12 2.71e-12 6.05e-12 6.26e-12
ERR 3.61e-12 1.83e-11 1.76e-11 7.87e-12
[…]
CPU 0.09 0.12 0.10 0.13
RES 1.91e-12 4.12e-12 4.42e-12 4.78e-12
ERR 3.02e-12 3.45e-11 1.28e-11 5.05e-12
[…]
CPU 0.11 0.14 0.12 0.15
RES 7.96e-12 8.98e-12 3.04e-12 3.07e-12
ERR 1.81e-11 1.44e-10 9.70e-12 5.23e-12

### v3

@@
Abstract. We present a stationary iteration based upon a block splitting for a class of indefinite least squares problem. Convergence of the proposed method is investigated and optimal value of the involving parameter is used. The induced preconditioner is applied to accelerate the convergence of the GMRES method for solving the problem. We also analysed the eigenpair distribution of the preconditioned matrix. We assess the efficiency of the proposed preconditioner by presenting results from a numerical comparison with several existing preconditioners.
Keywords: block splitting, indefinite least squares, preconditioning, convergence, GMRES, optimal parameter.
AMS Subject Classification: 65F10, 65F50, 65F08..
@@ 1 Introduction
$J=\begin{pmatrix}I_{p}&0\\ 0&-I_{q}\end{pmatrix},\qquad p+q=m.$ (2)
@@
Direct computations reveal that
$f(x):=(b-Ax)^{T}J(b-Ax)=x^{T}A^{T}JAx-2x^{T}A^{T}Jb+b^{T}Jb.$
So, we get
$\nabla f=2(A^{T}JAx-A^{T}Jb),$ (3)
and the corresponding normal equations can be written as
@@
It follows from (4) that unlike the least squares problem, the ILS problem does not necessarily have a solution. It follows from (3) that the Hessian matrix for the function $f$ is $H=2A^{T}JA$. So the minimization problem (1) has a unique solution if and only if the matrix $A^{T}JA$ is symmetric positive definite (SPD) (see [5, 8]). We assume that this condition holds throughout the paper.
When the problem size is small, Chandrasekaran et al. in [5] proposed a direct method based on the QR decomposition and Cholesky factorization for solving the problem. Later, Bojanczyk et al. [4] implemented a hyperbolic QR factorization of $A$ using Householder and hyperbolic Givens transformations. A backward stable hyperbolic QR factorization method for solving the ILS problem was presented by Xu in [19]. However, when the problem is large, the computational cost becomes prohibitively high. So it is recommended to use iteration methods for solving the problem.
In [11], Liu and Liu presented the two versions of the block SOR method for solving the problem. In [17], Song constructed an unsymmetric SOR (USSOR) method for solving the problem. A variable parameter Uzawa method for solving the indefinite least squares problem was presented by Meng et al. in [14]. Meng et al. in [13] proposed an enhanced Landweber method with momentum acceleration for solving large indefinite least squares problem. Recently, Meng et al. in [12] have stated alternating direction implicit iteration method for the problem.
@@
$A=\begin{pmatrix}A_{1}\cr A_{2}\end{pmatrix},\qquad b=\begin{pmatrix}b_{1}\cr b_{2}\end{pmatrix}$, (5)
@@
$\mathcal{B}$ $\hskip-5.69046pt=\hskip-5.69046pt$ $\begin{pmatrix}I&0&0\cr 0&P&0\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&-A_{1}&0\cr 0&0&-A_{2}^{T}\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{1}-\mathcal{N}_{1},$ (7)
$\mathcal{B}$ $\hskip-5.69046pt=\hskip-5.69046pt$ $\begin{pmatrix}I&0&0\cr 0&P&A_{2}^{T}\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&-A_{1}&0\cr 0&0&0\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{2}-\mathcal{N}_{2},$ (8)
$\mathcal{B}$ $\hskip-5.69046pt=\hskip-5.69046pt$ $\begin{pmatrix}I&A_{1}&0\cr 0&P&0\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&0&0\cr 0&0&-A_{2}^{T}\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{3}-\mathcal{N}_{3},$ (9)
and proved that the stationary iteration methods corresponding to the above splittings are unconditionally convergent. Recently, Li et al. in [10] split the matrix $\mathcal{B}$ as
$\mathcal{B}=\begin{pmatrix}I&A_{1}&0\cr 0&P&A_{2}^{T}\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&0&0\cr 0&0&0\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{4}-\mathcal{N}_{4},$ (10)
and showed that the corresponding stationary iteration method is also unconditionally convergent. The authors of [10, 18] applied the matrices $\mathcal{M}_{i}$ $(i=1,2,3,4)$ as preconditioners to accelerate the convergence of GMRES for solving the system (11). The numerical results presented in [10] demonstrate that the preconditioner $\mathcal{M}_{4}$ outperforms the preconditioners $\mathcal{M}_{i}$ $(i=1,2,3)$.
In this paper, we consider the following reformulation used in [17] of the problem (4)
where $I$ denotes the identity matrix of the appropriate dimension and $\delta=(\delta_{1};\delta_{2})=b-Ax$. By premultiplying the first row of the above system by $A_{1}^{T}$ and letting $\hat{\delta}_{1}=A^{T}\delta_{1}$ and $\hat{b}_{1}=A_{1}^{T}b$, we obtain the following system
where $P=A_{1}^{T}A_{1}$. In this case, all the block diagonal matrices are nonsingular. It is noted that the dimension of the coefficient matrix of this system is $\textbf{n}=2n+q$. We propose a parameterized block splitting (PBS) iteration method for the system (4), investigate convergence of the method and analyze the induced preconditioner.
@@ 2 The proposed iteration method
We now consider the stationary iteration method induced by the parametrized block splitting (PBS) $\mathcal{A}=\mathcal{M}_{\alpha}-\mathcal{N}_{\alpha}$, where
@@
$\begin{cases}x^{(k+1)}=P^{-1}(-\hat{\delta}_{1}^{(k)}+\hat{b}_{1}),\\ \delta_{2}^{(k+1)}=-\alpha A_{2}x^{(k+1)}+(\alpha-1)A_{2}x^{(k)}+b_{2},\qquad k=0,1,\ldots,\\ \hat{\delta}_{1}^{(k+1)}=A_{2}^{T}\delta_{2}^{(k+1)},\end{cases}$ (15)
@@
The iteration scheme (14) converges to the solution of the system (12) for any arbitrary initial guess if and only if $\rho(\mathcal{G}_{\alpha})<1$ [15]. In the sequel we investigate the convergence of the proposed method. We first recall two lemmas.
@@
Let $A\in\mathbb{R}^{m\times n}$ and the matrix $A^{T}JA$ be SPD, where the matrix $J$ was defined in Eq. (2). Suppose that the matrix $A_{1}$ defined in Eq. (5) is of full column rank. If $\mu$ is an eigenvalue of the matrix $Q=(A_{1}^{T}A_{1})^{-1}(A_{2}^{T}A_{2})$, then $0\leq\mu<1$.
@@
Considering that the matrix $A^{T}JA=A_{1}^{T}A_{1}-A_{2}^{T}A_{2}$ is SPD, it follows from Lemma 3.2 in [11] or Theorem 7.7.3 in [9]. ∎
@@
Let $(\lambda,\textbf{w})$ be an eigenpair of the iteration matrix $\mathcal{G}_{\alpha}$, i.e., $\mathcal{G}_{\alpha}\textbf{w}=\lambda\textbf{w}$. This means that $\mathcal{N}_{\alpha}\textbf{w}=\lambda\mathcal{M}_{\alpha}\textbf{w}$. So if we assume that $\textbf{w}=(x;y;z)$, then
@@
From Lemma 2, we have $0\leq\mu<1$. For $\mu=0$, both of the above inequalities hold true. So, we assume that $\mu\neq 0$. Clearly, the second equation holds true. From the first equation we see that
@@
$\alpha_{opt}={\rm arg}\hskip-17.07182pt\min_{\hskip-5.69046pt0<\alpha<1+\frac{1}{\mu_{\max}}}\rho(\mathcal{G}_{\alpha})=\frac{2}{1+\sqrt{1-\mu_{\max}}}.$
@@
We consider the equation (17) and without lose of generality we assume that $\mu\neq 0$. In this case, this equation has two roots
@@ 3 The PBS preconditioner
In the previous section, the PBS method is convergent if and only if $\alpha\in(0,1+1/\mu_{\max})$. It means that the eigenvalues of the matrix
$\mathcal{M}_{\alpha}^{-1}\mathcal{A}=\mathcal{I}-\mathcal{G}_{\alpha},$ (18)
@@
$\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}z_{1}\cr z_{2}\cr z_{3}\end{pmatrix}=\begin{pmatrix}r_{1}\cr r_{2}\cr r_{3}\end{pmatrix}$,
can be done via the following algorithm.
Algorithm 1. Solution of $\mathcal{M}_{\alpha}(z_{1};z_{2};z_{3})=(r_{1};r_{2};r_{3})$
@@
    Compute $z_{3}=r_{3}+A_{2}^{T}z_{2}$.
The coefficient matrix of the system in Step 1 is SPD. So, it can be solved directly using the Cholesky factorization for inexactly using the conjugate gradient method.
If the system of Step 1 is solved using the Cholesky factorization of the matrix $P$, then the computational complexity of the algorithm is $\mathcal{O}(n^{3})$, since the computational complexity of solving the system $Pz_{1}=r_{1}$ is $\mathcal{O}(n^{3})$ along with the computational complexity of matrix-vector operations in Steps 2 and 3 which is $\mathcal{O}(n^{2})$. Comparing the PBS preconditioner with the preconditioners $\mathcal{M}_{i}$ ($i=1,2,3,4$) reveals that for large systems, their computational complexities are almost the same.
The next theorem further analyses the eigenpairs distribution of the preconditioned matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$.
@@
    If $\eta\neq 1$ (for every $\alpha$), then the corresponding eigenvector is given by
    $\left(\frac{1}{\eta-1}P^{-1}A_{2}^{T}y;y;A_{2}^{T}y\right),$
@@
    If $\alpha=\eta=1$, then every vector of the form $(x;y;0)$ with at least one of the vectors $x$ and $y$ being nonzero is the corresponding eigenvector.
@@
    If $\alpha\neq 1$ and $\eta=1$, then every nonzero vector $(x;y;0)$ with $x\in\mathcal{N}(A_{2})$ is the corresponding eigenvector.
@@
Let $(\eta,\textbf{w})$ be an eigenpair of the matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$. It follows from Eq. (18) that each eigenvalue $\eta$ of $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$ can be expressed as $\eta=1-\lambda$, where $\lambda$ is an eigenvalue of $\mathcal{G}_{\alpha}$. Substituting $\lambda=1-\eta$ in Eq. (17) results in the following quadratic equation
$\eta^{2}+(\alpha\mu-2)\eta+1-\mu=0.$
So each eigenvalue $\eta$ of the preconditioned $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$ is a solution of this equation.
Letting $\textbf{w}=(x;y;z)$, we have $\mathcal{M}_{\alpha}^{-1}\mathcal{A}\textbf{w}=\eta\textbf{w}$, which can be written as
$\begin{pmatrix}P&0&I\cr A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr y\cr z\end{pmatrix}=\eta\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr y\cr z\end{pmatrix}$,
@@
$\begin{cases}(\eta-1)Px=z,\\ (1-\alpha\eta)A_{2}x=(\eta-1)y,\\ (\eta-1)(z-A_{2}^{T}y)=0.\end{cases}$ (20)
To prove $(i)$, we first show that $y\neq 0$. If $y=0$, then from the third equation in (20) we get $z=0$. Substituting $z=0$ in the first equation yields $x=0$. So, in this case we deduce that $\textbf{w}=(x;y;z)=0$ which is a contradiction with the vector w being an eigenvector. Now, from the third equation we see that $z=A_{2}^{T}y$. Substituting $z=A_{2}^{T}y$ in the first equation yields
$x=\frac{1}{\eta-1}P^{-1}z=\frac{1}{\eta-1}P^{-1}A_{2}^{T}y.$
To prove $(ii)$, if $\alpha=\eta=1$ then it follows from (20) that $z=0$. It is noted that for every vectors $x$ and $y$, the Eq. (20) holds true that proves $(ii)$.
To prove $(iii)$, we see that $z=0$ and $(1-\alpha)A_{2}x=0$ which implies $A_{2}x=0$, meaning that $x\in\mathcal{N}(A_{2})$. Hence the desired results is obtained. ∎
@@ 4 Numerical experiments
In this section, we present some numerical experiments to verify the theoretical results and show the effectiveness of the PBS preconditioner. Comparison with other recently proposed preconditioners are also given. All the numerical experiments have been conducted using some Matlab codes on a PC with a Pentium Dual-Core CPU, 4 gigabytes of RAM, and the Windows 7 operating system.
@@
$A^{T}JA=A_{1}^{T}A_{1}-A_{2}^{T}A_{2}=\begin{pmatrix}35&10&16\\ 10&11&19\\ 16&19&44\\ \end{pmatrix}$,
which is a SPD matrix. So all the conditions of Theorems 1 and 2 hold true. On the other hand, we have $\mu_{\max}=0.4976$. So, the method is convergent for any $\alpha$ with
$0<\alpha<1+\frac{1}{\mu_{\max}}=1+\frac{1}{0.4976}=3.009.$
@@
We have solved the system (12) using the PBS method (Eq. (16)). The initial guess is set to be a zeros vector and the iteration is stopped as soon as residual norm of the system is reduced by a factor of $10^{11}$. The number of iterations for the convergence and the CPU CPU processing time in second for the optimum value of $\alpha$ and some other values have been presented in Table 1. The numerical results confirm the theoretical results.
@@
In this example, we use four matrices TOLS340,TOLS1090, TOLS2000 and TOLS400 from Matrix Market (https://math.nist.gov/MatrixMarket/). The Tolosa matrices are encountered in the stability analysis of an aircraft flight model. The notable modes of this system are characterized by complex eigenvalues, with their imaginary parts situated within a specified frequency range. The objective is to determine the eigenvalues that possess the largest imaginary components. This problem has been studied at CERFACS (Centre Européen de Recherche et de Formation Avancée en Calcul Scientifique) in collaboration with the Aerospatiale Aircraft division. These matrices with their generic properties, including the size ($n$), the number of nonzero entries (nnznnz) and the condition number estimation (Cond), are presented in Table 2.
@@
matrix     $n$      nnznnz      Cond
TOLS 340 340 2196 $2.4e+05$
TOLS1090 1090 3546 $2.1e+06$
TOLS2000 2000 5184 $6.9e+06$
TOLS4000 4000 8784 $2.7e+07$
We set $A_{1}=$ TOLS 340, TOLS1090, TOLS2000 and TOLS4000. In this case, we have $p=n=340,1090,2000$ and $4000$. For each system we set $q=n$ and $A_{2}=6I_{n}$, where $I_{n}$ is the identity matrix of order $n$. The entries of the vectors $b_{1}$ and $b_{2}$ are set to be all ones. We solve the system (12) with restarted version of GMRES(10) in conjunction with the PBS preconditioner. Left-preconditioning is used. A zero vector is used as an initial guess and the iteration is stopped as soon as the residual of the original system is reduced by a factor of $10^{11}$. The parameter $\alpha$ in the PBS method is set to be $1$. We also compare the numerical results of the PBS preconditioner with the preconditioners induced by the block splittings (7), (8), (9), and (10) (denoted by BS1, BS2, BS3 and BS3) applied to the system (12). In the implementation of the preconditioners $\mathcal{M}_{i}$ $(i=1,2,3,4)$ within the GMRES method, we need to solve systems of the form $\mathcal{M}_{i}(z_{1};z_{2};z_{3})=(r_{1};r_{2};r_{3})$ which can be accomplished as follows (see [10, 18]).
Algorithm 2. Solution of $\mathcal{M}_{1}(z_{1};z_{2};z_{3})=(r_{1};r_{2};r_{3})$
  • 1.
    Set $z_{1}=r_{1}$;
  • 2.
    Solve $Pz_{2}=r_{2}$ for $z_{2}$;
  • 3.
    Set $z_{3}=r_{3}$.
Algorithm 3. Solution of $\mathcal{M}_{2}(z_{1};z_{2};z_{3})=(r_{1};r_{2};r_{3})$
  • 1.
    Set $z_{1}=r_{1}$;
  • 2.
    Set $z_{3}=r_{3}$;
  • 3.
    Solve $Pz_{2}=r_{2}-A_{2}^{T}z_{3}$ for $z_{2}$.
Algorithm 4. Solution of $\mathcal{M}_{3}(z_{1};z_{2};z_{3})=(r_{1};r_{2};r_{3})$
  • 1.
    Solve $Pz_{2}=r_{2}$ for $z_{2}$;
  • 2.
    Compute $z_{1}=r_{1}-A_{1}z_{2}$;
  • 3.
    Set $z_{3}=r_{3}$.
Algorithm 5. Solution of $\mathcal{M}_{4}(z_{1};z_{2};z_{3})=(r_{1};r_{2};r_{3})$
  • 1.
    Set $z_{3}=r_{3}$
  • 2.
    Solve $Pz_{2}=r_{2}-A_{2}^{T}z_{3}$ for $z_{2}$;
  • 3.
    Set $z_{1}=r_{1}-A_{1}z_{2}$.
Numerical results have been presented in Table 2. In this table the number of iterations (Iter), elapsed CPU time in second (CPU) ,
@@
where $\textbf{x}_{k}$ is the computed solution at iteration $k$ and $\textbf{x}^{*}$ is the exact solution to the system. To show the effectiveness of the preconditioners we also examined the restarted version of GMRES(10) method without preconditioning (No-Prec). A dagger ($\dagger$}) shows that the method has not converged in 1000 iterations. As the numerical results show all the preconditioners are effective. We also see that the PBS preconditioner is more efficient that the other preconditioners in terms of the number of iterations and the CPU time.
@@
   $m\times n$ PBS BS1 BS2 BS3 BS4 No-Prec
$640\times 340$ Iter 14 32 18 34 16 $\dagger$}
CPU 0.12 0.23 0.19 0.23 0.18
RES 9.13e-12 3.79e-12 3.52e-13 7.31e-12 3.21e-12
ERR 9.22e-12 3.07e-11 9.16e-13 9.06e-12 6.39e-12
$2180\times 1090$ Iter 14 32 16 34 15 $\dagger$}
CPU 0.15 0.21 0.18 0.23 0.16
RES 2.58e-12 2.71e-12 6.05e-12 6.26e-12 6.43e-12
ERR 3.61e-12 1.83e-11 1.76e-11 7.87e-12 2.57e-11
$4000\times 2000$ Iter 14 32 16 34 15 $\dagger$}
CPU 0.17 0.24 0.19 0.25 0.19
RES 1.91e-12 4.12e-12 4.42e-12 4.78e-12 4.5e-12
ERR 3.02e-12 3.45e-11 1.28e-11 5.05e-12 1.80e-11
$8000\times 4000$ Iter 13 30 16 34 15 $\dagger$}
CPU 0.21 0.34 0.25 0.40 0.24
RES 7.96e-12 8.98e-12 3.04e-12 3.07e-12 3.6e-12
ERR 1.81e-11 1.44e-10 9.70e-12 5.23e-12 1.29e-11